# CoBRA: A Universal Strategyproof Confirmation Protocol for Quorum-based Proof-of-Stake Blockchains

### v2

@@ Appendix A Rational attacks
(1): Assume the leader $L_{r}$ proposing the next block at some round $r$ (w.l.o.g.) is malicious. If conflicting blocks already exist, the claim holds. Otherwise, all correct validators initially agree on the same ledger $T$. A malicious leader partitions the correct validators into two subsets, $H_{1}$ and $H_{2}$, where $|H_{1}|=\lfloor h/2\rfloor$ and $|H_{2}|=\lceil h/2\rceil$. Before the bound $\Delta^{*}$ elapses, the leader sends a conflicting block $b_{1}$ to $H_{1}$ and $b_{2}$ to $H_{2}$. Rational validators sign both $b_{1}$ and $b_{2}$ because they can present inclusion proofs for $b_{1}$ and $b_{2}$ to disjoint client sets $C_{1}$ and $C_{2}$ in exchange for off-chain assets. Each conflicting block gathers $q$ signatures since $f+k+\lfloor h/2\rfloor\geq q$ and $f+k+\lceil h/2\rceil\geq q$. Therefore, there is a valid certificate for the conflicting blocks $b_{1}$ and $b_{2}$.
(2): Let us fix any honest validator $p_{1}\in H_{1}$ that witnesses the valid certificate $QC(b_{1})$ for block $b_{1}$ at time $t_{1}$ [^9] We use $p_{i}$’s local time as a reference, as each validator measures elapsed time using its own clock.. Similarly, let $p_{2}\in H_{2}$ be an honest validator that witnesses the valid certificate $QC(b_{2})$ for a conflicting block $b_{2}$ at time $t_{2}$, and without loss of generality, assume $t_{1}\leq t_{2}$.
First, it must hold that $t_{2}-t_{1}\leq\Delta^{*}$. That is because every validator in $H_{2}$ will have receive $QC(b_{1})$ at time $t_{1}+\Delta^{*}$ and would not vote for the conflicting $QC(b_{2})$. Moreover, each validator will have received and witnessed both certificates by time $t_{2}+\Delta^{*}\leq t_{1}+2\Delta^{*}$. Thus, any two conflicting blocks with valid certificates will be observed by all honest validators within $2\Delta^{*}$.
Liveness Attacks
Censoring transactions. In traditional Proof-of-Stake blockchains, the majority of transaction fees are awarded to the validator that successfully proposes a block, incentivizing the inclusion of as many transactions as possible. However, if a proposer fails to produce a valid block, the pool of unprocessed transactions grows, potentially increasing the rewards available to subsequent proposers. This dynamic creates an incentive for rational validators to launch censorship attacks [35] against competing proposers to maximize their own rewards.
Bribery attacks. Byzantine validators can use their share of the rewards to bribe rational validators and censor all correct validators. As an example, we consider a $12f+$ 1–commitable SMR protocol with $f$ Byzantine, $f$ rational, and $f+1$ correct validators. To successfully launch the attack, Byzantine validators only need to bribe a single rational validator $p$, by distributing all their rewards to $p$. In that scenario, all the rest block proposers gather at most $2f$ signatures for their proposals. Consequently, only the byzantine validators along with $p$ would successfully propose a block. As a result, the adversary would manage to censor any transaction held by correct validators.
Appendix B Omitted proofs
Lemma 9.
Assume that correct validators employ the Strong Chain finalization rule and at least one coalition of validators forks the system. Any rational validator participating in the attack can double spend at most $D$ coins.
Proof.
Similar to Lemma 2, we will upper bound the total amount of coins in conflicting $2\Delta^{*}$-recent blocks finalized by at least one validator.
Case 1: Every correct validator finalizes up to $D$. The proof is similar to Lemma 2.
Case 2: At least one correct validator $p_{i}$ monitors an $i$-strong chain $L_{p_{i}}$ with $i>f/4$. Since $1+i$ correct validators monitor $L_{p_{i}}$, there can be at most $f-i$ additional conflicting ledgers. We assume the worst case, which is that $f-i$ conflicting ledgers exist. We fix any rational validator $p$ engaging in the attack, participating in $f-m-i$ conflicting ledgers $L_{i\in[f-m-i]}$, with $m\in[f-i]$.
For each $i\in[f-m-i]$, at most $m+1$ correct validators and therefore at least $f-m$ rational validators monitor ledger $L_{i}$. We prove that by contradiction. Assume there is a ledger $L_{i}$ with $m+2$ correct validators. At least $i+1$ correct validators monitor $L_{p_{i}}$, at least $m+2$ correct validators monitor $L_{i}$ and at least one correct validator exists in the rest $f-m-i-1$ conflicting ledgers $L_{j\neq i}$. In total, this accounts for $f+2$ correct validators, leading to a contradiction. Furthermore, since there are at most $m+1$ correct validators per ledger, the total amount of coins finalized by correct validators on any ledger is at most $\frac{f}{f-m}D$.
Finally, there are $f-m-i$ conflicting ledgers with at most $f-m$ rational validators spending up to $\frac{f}{f-m}D$ per ledger. So, the amount of double-spent coins is bounded by $max_{\{m\in[f-i]\}}\frac{(f-m-i)f}{(f-m)^{2}}D$ which is bounded by $D$ for $i\in\{\frac{f}{4},...,f\}$. We can show that, by taking the continuous extension of the function $F(m)=\frac{(f-m-i)f}{(f-m)^{2}}D$ for $m\in[0,f-i]$ and some constant $i\in\{\lfloor\frac{f}{4}+1\rfloor,...,f\}$, which takes its maximum value for $m=f-2i$. Since the function takes maximum value at an integer value, the maximum will be the same for the discrete function and thus $\frac{f}{4i}D\leq D$.
Lemma 10.
At any round, there can only exist a unique strongest chain.
Proof.
For the sake of contradiction, assume that at some round $r$, there exist two conflicting chains $T_{1}$, $T_{2}$ that both form a strongest chain. That means that $T_{1}$ is an $i_{1}$-strong chain for some $i_{1}>\frac{f-1}{2}$ and $T_{2}$ is an $i_{2}$-strong chain for some $i_{2}>\frac{f-1}{2}$. Therefore, there exist two distinct subsets $H_{1}$ and $H_{2}$ that consist of $1+i_{1}$ correct validators participating in $T_{1}$ and $1+i_{2}$ correct validators participating in $T_{2}$ respectively. However, $11+i_{1}+1+i_{2}>f+$ 1. Since there are $f+1$ only correct validators in the system, we reached a contradiction. ∎
Lemma 11.
Assume that correct validators employ the Strongest Chain finalization rule and at least one coalition of validators forks the system. Any rational validator participating in the attack can double-spend at most $D$ coins.
Proof.
We will prove that when validators employ the strongest chain finalization rule, rational validators participating in coalitions can double-spend up to $D$ coins. Similar to Lemma 2, we will upper bound the total amount of coins spent in conflicting $2\Delta^{*}$-recent finalized blocks.
Case 1: There is no correct validator $p_{i}$ monitoring an $i_{1}$-strong chain with $i_{1}>(f+1)/2$. If every correct validator finalizes up to $D$, the proof follows from Lemma 2. If at least one correct validator monitors an $i_{1}$-strong chain with $i_{1}>f/4$, the proof follows from Lemma 9.
Case 2: At least one correct validator $p_{i}$ monitors an $i_{1}$-strong chain with $i_{1}>(f+1)/2$. By upper bounding the total coin value in the last $2\Delta^{*}$ blocks of conflicting ledgers finalized by at least one validator, we also bound the double-spent coins per misbehaving validator.
Case 2a: Every correct validator in a conflicting ledger, finalizes up to $D$. A number of $1+i_{1}$ correct validators monitor $L_{p_{i}}$, and thus there are up to $f-i_{1}$ extra conflicting ledgers. Assume that there are $f-i_{1}$ extra conflicting ledgers and let us any rational validator participating in $f-m-i_{1}$ of the conflicting ledgers for $m\in[f-i]$. In each fork, there are at most $m+1$ correct validators and at least $f-m$ rational validators (with the same analysis done in Lemma 2). Therefore, rational misbehaving validators can double-spend up to $max_{\{m\in[f-i_{1}]\}}\frac{(f-m-i_{1})}{(f-m)}D\leq{m\in[f-i_{1}]\}}\frac{(f-m)}{(f-m)}D=D$.
Now, we fix any rational validator $p$ participating in $L_{p_{i}}$ and in $i_{3}-m$ of the conflicting ledgers $L_{i\in[i_{3}-m]}$ for $m\in[i_{3}]$. We will upper bound the coins that $p$ can spend in these conflicting blocks. First, in $L_{p_{j}}$, there are finalized up to $\frac{f}{(f-m)(f-i_{2})}D<A(m)=\frac{f}{(f-m)(f-\frac{f-1}{2})}D$. For the conflicting ledgers $L_{i\in[i_{3}-m]}$, we can show that there are at most $m+1$ correct validators and at least $f-m$ rational validators per fork, with the same analysis done in 2. Thus, each validator can double-spend up to $\frac{f+1-i_{1}^{*}-i_{2}^{*}-m}{f-m}D=(1+\frac{1-i_{1}^{*}-i_{2}^{*}}{f-m})D\leq B(m)=1-\frac{3(f+2)}{4(f-m)})D$. Now we need to upper bound $C(m)=A(m)+B(m)=(1-\frac{3f^{2}+f+6}{4(f-m)(f+1)})D<D$ since $m\in[f+1-i_{1}^{*}-i_{2}^{*}]$ and $f>0$.
@@ Appendix C Impossibility results
For the sake of contradiction assume that there is a q–commitable SMR protocol $\Pi$ that is $(n,k,f)$-resilient for $f\geq\lceil h/2\rceil$, where $h=n-k-f$ is the number of correct validators. Assume that each validator has deposited $D\in\mathbf{N}$ stake to participate in $\Pi$ and assume any period $\Delta_{W}$ where a validator can withdraw their stake.
We will construct a run where colluding with Byzantine validators to create valid certificates for conflicting blocks is a dominant strategy for rational validators. This will lead to a safety violation.
Double spending attack. Since $f\geq\lceil h/2\rceil$, it holds that $f+k+\lfloor h/2\rfloor\geq q$. In Lemma 8, we showed that a malicious validator can partition correct validators in two distinct subsets $H_{1}$ and $H_{2}$ monitoring two conflicting ledgers $T_{c}\cup T_{1}$ and $T_{c}\cup T_{2}$ respectively, where $T_{c}$ is the common part of the ledgers. Assume the attack takes place during an interval  [^10] We can use a reference point the local clock of any correct validator. $[t,t^{\prime}]$ where $t^{\prime}>t+\Delta_{W}$ and let $\mathcal{F}$ be the set of the $f+k$ misbehaving validators.
@@
World 1. (GST = 0.) Validators in $H_{1}$ and $F$ are correct. Validators in $H_{2}$ are Byzantine and do not participate in the protocol. The input to $\Pi$ are the transactions in $Tx_{1}$. Validator $p_{1}\in H_{1}$ witnesses blocks in $T_{1}$  [^13] We mean that $p_{1}$ sequentially finalizes blocks in $T_{1}$. with quorum certificates of $n-f$ votes.
World 2. (GST = 0) This is symmetrical to World 1. Now, validators in $H_{2}$ and $F$ are correct. Validators in $H_{1}$ are Byzantine and do not participate in the protocol. The input to $\Pi$ are the transactions in $Tx_{2}$. Validator $p_{2}\in H_{2}$ witnesses blocks in $T_{2}$ along with the quorum certificates of $n-f$ votes.
@@
For the sake of contradiction assume that there is a q–commitable SMR protocol $\Pi$ that is $(n,k,f)$-resilient for $f\geq h$, where $h=n-k-f$ is the number of correct validators. Assume that each validator has deposited $D\in\mathbf{N}$ stake to participate in $\Pi$.
@@
Double spending attack. The input of $\Pi$ at time $t$ is a set of transactions $Tx=\cup_{i=1}^{C_{t}}Tx_{i}$ constructed as follows. There is a set $\mathcal{R}$ of at least $r=max(q-f,0)$ rational validators such that every validator in $\mathcal{R}$ transfers $c>0$ coins to the client $c_{i\in\{1,\dots,C_{t}\}}$ in $Tx_{i\in\{1,\dots,C_{t}\}}$. Moreover, we construct the run such that $(C_{t}-1)C>D$. Since the deposit $D$ is decided before the run of the protocol and $C_{t}$ is arbitrary, it is feasible to construct such a run.
@@
Client $c_{i}$ cannot distinguish between World 1 and World 2 and therefore, the rule $\rho_{1}$ is not sufficient. Now, consider any rule $\rho_{1}$ which requires stages of at least $t$ votes. Similar to the proof of Theorem 3, an inclusion proof of stages of $t=n-f$ votes is the strongest certificate that clients can receive. Therefore, each client $c_{i\in\{1,\dots,C_{t}\}}$ clients will accept the inclusion proofs for the respective block $b_{i}$.
@@
Assume, for contradiction, that there exists a $q$-participation SMR protocol $\Pi$ that is $(n,k,f)$-resilient for $f\geq\lceil h/2\rceil$, where $h=n-k-f$ represents the number of correct validators, and $\Pi$ does not incur a finality delay of at least $\Delta^{*}$. Moreover, suppose each validator has staked $D\in\mathbf{N}$ to take part in $\Pi$.
We will construct an attack where it is dominant strategy for rational validators to collude with Byzantine validators and produce valid certificates for conflicting blocks, leading to safety violation. Similar to the proof of Theorem 3, a malicious validator can partition correct validators in two distinct subsets $H_{1}$ and $H_{2}$ which monitor the conflicting ledgers $T_{c}\cup T_{1}$ and $T_{c}\cup T_{2}$ respectively. Additionally, there is a partition of the clients in two sets $C_{1},C_{2}$, and the input to the blockchain protocol is $Tx=Tx_{1}\cup Tx_{2}$, where a set $\mathcal{R}$ of at least $p-f-min(|H_{1}|,|H_{2}|)$ rational validators spend $D+\epsilon_{1}$ coins in $Tx_{1}$ to a subset of clients in $C_{1}$ and $D+\epsilon_{2}$ coins in $Tx_{2}$ to a subset of clients $C_{2}$ where $\epsilon_{1},\epsilon_{2}>0$. Then, every transaction in $Tx_{1}$ must be committed to $T_{1}$, and every transaction in $Tx_{2}$ must be committed to $T_{2}$ within $\Delta^{*}$ by assumption.
We will now show that such a run is feasible. Specifically, we will construct a run where, for any $D$, every rational validator spend $D+\epsilon_{1}$ coins in $T_{1}$ (same for $T_{2}$). To this end, cosnider the following notation. $G^{p}_{0}$ denotes the coins of validator $p$ in $T_{c}$. Then, we denote by $S^{p}_{i}$ the total coins paid to clients in $C_{1}$ by validator $p$ and by $R^{p}_{i}$ the participation rewards $p$ has received until the $i$-th block of $T_{1}$ (excluding $T_{c}$). We need to show that for every rational validator $p$ there is a $k$ s.t. $S^{p}_{k}>D$, or $G^{p}_{0}+R^{p}_{k}>D$.
We will show that the participation rewards $p$ receives cannot be bounded. Toward this direction, consider the first $i\in\mathcal{N}$ such that $R^{p}_{i}=\mathcal{C}^{p}_{0}$ as follows. Every validator in $\mathcal{R}$ spends all their coins, and therefore all coins belong to the hands of all the other users. When enough blocks has been processed in the system, any rational validator $p$ actively participating (by proposing or voting blocks depending on the reward scheme) will receive at least $\mathcal{C}^{p}_{0}$ in participation rewards. Specifically, w.l.o.g., if the participation rewards per block are $r$,$n$ blocks have been processed in the system and $p$ has participated in a percentage of $\alpha$ of them, then there is a run where $\alpha\cdot n\cdot r\geq\mathcal{C}^{p}_{0}$, which holds since it can be $n\to\infty$. Then, rational validators will spend all of their money to pay clients in $C_{1}$ and the above scenario is repeated for $j$ times such that there is $i\in\mathcal{N}$, $R^{p}_{i}=j\mathcal{C}^{p}_{0}$ and $(j+1)\mathcal{C}^{p}_{0}>D$. Since it can be that $j\to\infty$, this is a feasible scenario.
Theorems 3 and 3 share the same Byzantine resilience and assume similar network conditions during a partition lasting up to $\Delta^{*}$. Thus, the proof of Theorem 3 extends to show that: (1) safety is violated, with correct validators $p_{1}\in H_{1}$ and $p_{2}\in H_{2}$ finalizing conflicting blocks in ledgers $T_{1}$ and $T_{2}$, and (2) clients accept inclusion proofs backed by valid certificates. ∎

### v3

@@ B.3. Final Theorems
Assume an equivocation occurs, i.e., conflicting blocks containing finality votes are formed, and let time tt [^13] We fix the local clock fo any participant, as only elapsed time matters. be the first time at which finality certificates for conflicting blocks are formed.
Clients accept a block $b$ only if it carries a valid finality certificate $\mathcal{F}(b)$. Finality votes forming $\mathcal{F}(b)$ are included only in a subset of blocks $\mathcal{B}$ where each block in $\mathcal{B}$ carries a valid quorum certificate. Since $\mathcal{F}(b)$ contains $2f+1$ finality votes and honest validators issue finality votes only for blocks whose certificates satisfy the finalization rule, at least one honest validator must have witnessed a valid certificate for each conflicting block and broadcast the corresponding certificate. Therefore, at time $t+\Delta^{*}$ every honest validator will have witnessed the first conflicting certificates votes and stop finalizing new blocks (line 17 or 29). Thus, the last conflicting blocks must have been created before $t+\Delta^{*}$ and will be witnessed by all validatots by time $t+2\Delta^{*}$.
We now argue that all honest validators will witness the same set of conflicting finalized ledgers $\mathcal{L}$. Every block whose finality votes are included in valid quorum certificates observable by at least one honest validator will eventually be observed by all honest validators, since this honest validator broadcasts the certificate. We now consider the case where quorum certificates are initially withheld by Byzantine validators. Since the $2f+1$ finality votes forming a finality certificate are included in a block with an ss–phase quorum certificate, and honest validators vote for an ss–phase certificate only after witnessing a complete (s1)(s\!-\!1)–phase certificate for the same block, at least one honest validator must have observed the corresponding (s1)(s\!-\!1)–phase certificate for every finalized block and broadcast it (line 52). Consequently, even if some ss–phase certificates were initially withheld, within $2\Delta^{*}$ all honest validators receive the complete set of finalized blocks accepted by clients and construct the same set of conflicting ledgers $\mathcal{L}$.
Each validator returns to the state of its last $2\Delta^{*}$-finalized block; by Lemma B.1, no block conflicting with this block exists. After witnessing an equivocation, validators then wait at least $2\Delta^{*}$ before executing transactions. Consequently, each validator update their state only at some time $t^{\prime}\in[t+2\Delta^{*},t+3\Delta^{*}]$, by which time they have observed all conflicting blocks with finality certificates in $\mathcal{L}$. Using the same deterministic rule (e.g., ordering by the hash of the last finalized block), all validators execute transactions from the conflicting ledgers in the same order.
Any transaction issued by an honest sender is executed successfully, since the sender holds the corresponding coins in the respective ledger. Transactions whose recipients lack sufficient balance correspond to double-spent coins. By Lemmas B.2B.4B.6, within the interval $t,\,t+\Delta^{*}]$, at most $fD$ coins are double-spent in total. Because, all honest validators identify (the same) at least $f+1$ misbehaving validators, these coins are covered by the stake of misbehaving validators which is in total $(f+1)D>fD$; any remaining stake of misbehaving validators is burned (line 63).
Thus, all honest validators compute the same resulting state and sign the same state commitment by time $t+3\Delta^{*}$. In particular, at least $f+1$ honest validators sign the state commitment which they broadcast to every validator (line 64). Thus every honest validator will have received the verifiable state commitment by time $4\Delta^{*}$. In the resulting state, all affected clients are reimbursed and all misbehaving validators are penalized.
@@ Appendix C Impossibility results
C.1. Proof Sketches
Proof sketch of Theorem 3. The adversary can partition correct validators in two distinct subsets $H_{1}$, $H_{2}$ and perform the double-spending attack described in Lemma A.1.This attack can be beneficial for rational validators since for any value of stake $D$ they lock before the run, there is an input of conflicting transactions $T_{1}$ and $T_{2}$ where rational validators transfer all their liquid coins to distinct subset of clients $C_{1}$ and $C_{2}$ respectively. Notably, misbehaving validators can withdraw their stake without even being slashed.
Proof sketch of Theorem 3. We modify the attack described in Theorem 3 as follows. For any stake value $D$ locked by rational validators before the execution, there exists an input consisting of conflicting transactions $T_{1}$ and $T_{2}$ such that rational validators transfer a total of $qD$ coins to clients in $C_{1}$ and $C_{2}$, respectively, and hence double-spend $qD$ coins. Even if the attack is eventually detected and the misbehaving validators are slashed and forfeit their stake, or even if the stake of misbehaving validators is instead offered as a reward for reporting misbehavior, the attack remains strictly beneficial and therefore preferable for rational validators.
Such a configuration can be constructed whenever at least one of the following resources is unbounded: (i) the initial liquid coins held by rational validators, (ii) the coins available as bribes from malicious clients, or (iii) the participation rewards. In particular, when participation rewards are unbounded, rational validators can repeatedly reclaim their initial funds through participation rewards, spend them again in conflicting transactions, and repeat this process sufficiently many times before GST, thereby accumulating a total double-spent value exceeding $qD$. ∎
Proof sketch of Theorem 3. When $f+k\geq q$, Byzantine and rational validators can successfully construct valid certificates for an arbitrary number of conflicting blocks and provide them to clients. For any stake $D$ that rational validators have locked before the run, there is a configuration in which each rational validator can double-spend more than $D$ coins. That is because the number of clients is arbitrary and unknown before the run. Even if the conflicting blocks are detected and the misbehaving validators get slashed, the attack is still beneficial for rational validators. ∎
Proof sketch of Theorem 3. The adversary can perform the double-spend attacks described in Theorems 33, before the synchronous bound $\Delta^{*}$ elapses. That is because both cases assume the same resilience against Byzantine adversaries and operate under similar network conditions during a network partition in $\Delta^{*}$. ∎
C.2. Full proofs
@@
For the sake of contradiction assume that there is a q–commitable SMR protocol $\Pi$ that is $(n,k,f)$-resilient for $f\geq\lceil h/2\rceil$, where h=nkfh=n-k-f is the number of correct validators. Assume that each validator has deposited $D\in\mathbf{N}$ stake to participate in $\Pi$ and assume any period $\Delta_{W}$ where a validator can withdraw their stake.
We will construct a run where colluding with Byzantine validators to create valid certificates for conflicting blocks yields greater utility for rational validators.
Double spending attack. Since $f\geq\lceil h/2\rceil$, it holds that $f+k+\lfloor h/2\rfloor\geq q$. In Lemma A.1, we showed that a malicious validator can partition correct validators in two distinct subsets $H_{1}$ and $H_{2}$ monitoring two conflicting ledgers $T_{c}\cup T_{1}$ and $T_{c}\cup T_{2}$ respectively, where $T_{c}$ is the common part of the ledgers. Assume the attack takes place during an interval  [^14] We can use a reference point the local clock of any correct validator. $t,t^{\prime}]$ where $t^{\prime}>t+\Delta_{W}$ and let $\mathcal{F}$ be the set of the $f+k$ misbehaving validators.
@@
World 1. (GST = 0.) Validators in $H_{1}$ and $F$ are correct. Validators in $H_{2}$ are Byzantine and do not participate in the protocol. The input to $\Pi$ are the transactions in $Tx_{1}$. Validator $p_{1}\in H_{1}$ witnesses blocks in $T_{1}$  [^17] We mean that $p_{1}$ sequentially finalizes blocks in $T_{1}$. with quorum certificates of nfn-f votes.
World 2. (GST = 0) This is symmetrical to World 1. Now, validators in $H_{2}$ and $F$ are correct. Validators in $H_{1}$ are Byzantine and do not participate in the protocol. The input to $\Pi$ are the transactions in $Tx_{2}$. Validator $p_{2}\in H_{2}$ witnesses blocks in $T_{2}$ along with the quorum certificates of nfn-f votes.
@@
For the sake of contradiction, assume that there exists a qq–commitable SMR protocol $\Pi$ that is $(n,k,f)$-resilient for some $q$ s.t. $q\leq f+k+\lfloor h/2\rfloor$, where h=nkfh=n-k-f is the number of correct validators. Also, assume that each validator has deposited $D\in\mathbf{N}$ stake to participate in $\Pi$.
We construct an attack in which rational validators collude with Byzantine validators to produce valid certificates for conflicting blocks, yielding strictly greater utility for rational validators and leading to a safety violation. Similar to the proof of Theorem 3, a malicious validator can partition correct validators in two distinct subsets $H_{1}$ and $H_{2}$ which monitor the conflicting ledgers $T_{c}\cup T_{1}$ and $T_{c}\cup T_{2}$ respectively. Additionally, there is a partition of the clients in two sets $C_{1},C_{2}$, and the input to the blockchain protocol is $Tx=Tx_{1}\cup Tx_{2}$, where a set $\mathcal{R}$ of at least $p-f-min(|H_{1}|,|H_{2}|)$ rational validators spend $qD$ coins in $Tx_{1}$ to a subset of clients in $C_{1}$ and $qD$ coins in $Tx_{2}$ to a subset of clients $C_{2}$ where $\epsilon_{1},\epsilon_{2}>0$. Therefore, even if the attack is eventually detected and rational validators forfeit their stake, or if the stake of misbehaving validators is instead offered as a reward for reporting misbehavior, the total value double-spent makes the attack strictly beneficial for rational validators.
We will now construct a run where, for any $D$, every rational validator spend $qD$ coins in $T_{1}$ (same for $T_{2}$). To this end, cosnider the following notation. $G^{p}_{0}$ denotes the coins of validator $p$ in $T_{c}$ plus the maximum amount of bribes from malicious validators to fork the system. Then, we denote by $S^{p}_{i}$ the total coins paid to clients in $C_{1}$ by validator $p$ and by $R^{p}_{i}$ the participation rewards $p$ has received until the $i$-th block of $T_{1}$ (excluding $T_{c}$).
We need to show that for every rational validator $p$ there is a $k$ s.t. $S^{p}_{k}>qD$, or $G^{p}_{0}+R^{p}_{k}>qD$. If the initial coins held by $p$, or the coins provided by malicious clients as bribes, are unbounded, this condition trivially holds. We now show that the condition also holds if the participation rewards $R^{p}_{k}$ are unbounded.
Toward this direction, consider the first $i\in\mathcal{N}$ such that $R^{p}_{i}=\mathcal{C}^{p}_{0}$ as follows. Every validator in $\mathcal{R}$ spends all their coins, and therefore all coins belong to the hands of all the other users. When enough blocks has been processed in the system, any rational validator $p$ actively participating (by proposing or voting blocks depending on the reward scheme) will receive at least $\mathcal{C}^{p}_{0}$ in participation rewards. Specifically, w.l.o.g., if the participation rewards per block are $r$,$n$ blocks have been processed in the system and $p$ has participated in a percentage of $\alpha$ of them, then there is a run where $\alpha\cdot n\cdot r\geq\mathcal{C}^{p}_{0}$, which holds since it can be $n\to\infty$. Then, rational validators will spend all of their money to pay clients in $C_{1}$ and the above scenario is repeated for $j$ times such that there is $i\in\mathcal{N}$, $R^{p}_{i}=j\mathcal{C}^{p}_{0}$ and $(j+1)\mathcal{C}^{p}_{0}>qD$. Since it can be that $j\to\infty$, this is a feasible scenario. ∎
Proof of Theorem 3
Proof.
For the sake of contradiction assume that there is a q–commitable SMR protocol $\Pi$ that is $(n,k,f)$-resilient for $f\geq h$, where h=nkfh=n-k-f is the number of correct validators. Assume that each validator has deposited $D\in\mathbf{N}$ stake to participate in $\Pi$.
@@
Double spending attack. The input of $\Pi$ at time $t$ is a set of transactions $Tx=\cup_{i=1}^{C_{t}}Tx_{i}$ constructed as follows. There is a set $\mathcal{R}$ of at least r=max(qf,0)r=max(q-f,0) rational validators such that every validator in $\mathcal{R}$ transfers $c>0$ coins to the client $c_{i\in\{1,\dots,C_{t}\}}$ in $Tx_{i\in\{1,\dots,C_{t}\}}$. Moreover, we construct the run such that $(C_{t}-1)C>D$. Since the deposit $D$ is decided before the run of the protocol and $C_{t}$ is arbitrary, it is feasible to construct such a run.
@@
Client $c_{i}$ cannot distinguish between World 1 and World 2 and therefore, the rule $\rho_{1}$ is not sufficient. Now, consider any rule $\rho_{1}$ which requires stages of at least $t$ votes. Similar to the proof of Theorem 3, an inclusion proof of stages of t=nft=n-f votes is the strongest certificate that clients can receive. Therefore, each client $c_{i\in\{1,\dots,C_{t}\}}$ clients will accept the inclusion proofs for the respective block $b_{i}$.
@@
Proof.
Assume, for contradiction, that there exists a qq–commitable SMR protocol $\Pi$ that is $(n,k,f)$-resilient for i) $f\geq\lceil h/2\rceil$, ii) or , where h=nkfh=n-k-f represents the number of correct validators, and $\Pi$ does not incur a finality delay of at least $\Delta^{*}$. Moreover, suppose each validator has staked $D\in\mathbf{N}$ to take part in $\Pi$.
Theorems 3 and 3 share the same Byzantine resilience and assume similar network conditions during a partition lasting up to $\Delta^{*}$, as in Theorem 3, with $f\geq\lceil h/2\rceil$ and $q\leq f+k+\lfloor h/2\rfloor$, respectively. Thus, the proofs of Theorems 3 and 3 extend to show that: (1) safety is violated, with correct validators $p_{1}\in H_{1}$ and $p_{2}\in H_{2}$ finalizing conflicting blocks in ledgers $T_{1}$ and $T_{2}$, and (2) clients accept inclusion proofs backed by valid certificates. ∎